crispyrat:

If x is a real number, find 49x^2+14x(19-7x)+(19-7x)^2. so so far i tried to simplfy this eqaution to 49x^2+(19-7x)(14x+19-7x)=49x^2+(19-7x)(7x+19) but idk what next ;-;

AZ:

no, it's similar to that question you asked previously you have to use the formula \( a^2 + 2ab + b^2 = (a+b)^2\) so your question is \( 49x^2 + 14x(19-7x) + \color{orange}{(19-7x)}^2\) do you see how you can write 49x^2 as (7x)^2 \( \color{red}{(7x)}^2 + 14x(19-7x) + \color{orange}{(19-7x)}^2\)

crispyrat:

i see it but i am a bit confused on what to do on that. can you explain in a different way?

AZ:

well your question is a bit incomplete? is that whole equation equal to 0?

crispyrat:

i guess thats all i have sadly :(

1 attachment
AZ:

but otherwise basically \( \color{red}{a}^2 + 2\color{red}{a}\color{orange}{b} + \color{orange}{b}^2 = (\color{red}{a}+\color{orange}{b})^2\) \( \color{red}{(7x)}^2 + 14x(19-7x) + \color{orange}{(19-7x)}^2 = ??\) and so you can just factor the equation

crispyrat:

oh so (7x+19-7x)^2 and then is x (19)(19)???

AZ:

yeah (7x+19-7x)^2 is correct I guess they just wanted you to factor it?

crispyrat:

so x is 361?

AZ:

Yup! because 7x - 7x = 0 and you're left with 19^2

AZ:

I didn't realize immediately that 7x would cancel out with -7x haha

crispyrat:

thx so much :)

AZ:

you're welcome :)